OpenStudy (anonymous):

The centroid of equilateral triangle is (0,0) and length of altitude is 6.The equation of circumcircle of triangle is ?

OpenStudy (shiraz14):

Equation of circumcircle: x^2 + y^2 = 16

OpenStudy (kc_kennylau):

1. Find the radius of the circumcircle 2. Substitute the center and the radius into the general equation \((x-a)^2+(y-b)^2=r^2\), where the center is \((a,b)\) and the radius is \(r\)